OpenStudy (anonymous):

Which of the following is the equation of the axis of symmetry of the quadratic equation y = 2x2 + 24x + 62? answers please

OpenStudy (anonymous):

This one is in standard form. To find the vertex we can use -b/2a = -24/4 = -6 The axis of symmetry is given by x= -6

OpenStudy (slaaibak):

First step is to write it in this form: y = a(x+b)^2 + c where x=-b is the axis of symmetry. To obtain that form, factor and use completing of the square

OpenStudy (slaaibak):

^Or that

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

or you can find the two x-intercepts and then find the midpoint of the two